# Bayesian Inference of Gravity through Realistic 3D Modeling of Wide Binary Orbits: General Algorithm and a Pilot Study with HARPS Radial Velocities

### v4

@@ 3.2 Bayesian Inference of the Gravitational Anomaly Parameter $\Gamma$
While Newtonian gravity requires Binary #24 to be a fly-by or chance-association in the 3D space, it is not so, based on MOND gravity theories, because $v/v_{\rm esc}<1$ (see Table 2) is feasible in MOND (existing MOND gravity theories (Bekenstein:1984; Milgrom:2010) predict a velocity boost of about 20%). Thus, if the scientific question is to test standard and nonstandard theories with wide binaries, #24 could be taken as evidence for nonstandard theories. It is also important to remember that this system survived stringent observational selection criteria just like all the other systems carefully collected by Saglia:2025. Moreover, the random chance-association probability for this system is extremely low or essentially zero as can be estimated through a Monte Carlo method. Specifically, the expected number of chance-association systems within 100 pc (like #24) for the sky-projected separation <6<6 kau, the line-of-sight separation <0.5<0.5 pc (considering $2\sigma$ uncertainties in the distance measurements), and the 3D relative speed <1<1 km s-1, is $\lesssim 8\times 10^{-4}$ for the one-dimensional velocity dispersion >30>30 km s-1 of peculiar velocities which is a conservative limit for solar-type stars in the solar neighborhood. Thus, it is extremely unlike that one such system was observed by chance out of a limited database which does not even include all stars within 100 pc.
@@
Figure 9 explicitly compares the observed 3D velocity components and the total mass of Binary #24 with the best-fit (MLE) values predicted by Newtonian, MOND-type ($G=1.4G_{\rm N}$), and generalized gravity ($\Gamma$ free) models. Newtonian model shows a large discrepancy in $v_{z^{\prime}}$ and a moderate discrepancy in $\log_{10}f_{M}$ while the other models reproduce the observed quantities well. The difference between the Newtonian prediction and the observed value of $v_{z^{\prime}}$ is $111$ m s-1 which is $7.6$ times the nominal measurement uncertainty and $2.6$ times the enlarged uncertainty considering possible errors in correcting for the effects of gravitational redshift and convective flow (see Saglia:2025). Based on the enlarged error of $v_{z^{\prime}}$, $\Delta{\rm BIC}$ (see Appendix A) between the Newtonian model and the other models is $\approx 8$ which is very significant. Moreover, the Newtonian model predicts an extremely unlikely (or impossible) combination of $(e,\Delta\phi)=(1.00,338.2^{\circ})$. It is interesting to note that the pseudo-Newtonian model with $G=1.4G_{\rm N}$ predicts $(e,\Delta\phi)=(0.91,339.7^{\circ})$, which seems somewhat rare but not impossible, while the generalized gravity model predicts $(e,\Delta\phi)=(0.67,203.6^{\circ})$. Because $z^{\prime}$ is not used in obtaining the MLE (see Appendix A), no models match the observed $z^{\prime}$ (nearly) exactly. Interestingly, all models predict $z^{\prime}\approx 0$, which is within its (large) measurement uncertainty.
@@
Although it is extremely unlikely that Binary #24 is a chance association, I also consider results without it just in case the system may be kinematically contaminated for unknown reasons. The thick dotted red curves in the left panels of Figure 8 show the consolidated PDFs of $\Gamma$: $\Gamma=0.063_{-0.041}^{+0.065}$ (thermal prior) or $0.051_{-0.039}^{+0.045}$ (flat prior). These results are only weakly discrepant with Newton and cannot distinguish well between Newton and MOND. This is not surprising given the small number of wide binaries. What is surprising is the presence of Binary #24 whose inclusion results in a discrepancy with Newton. It will be interesting to see whether future observational studies can discover more wide binaries like Binary #24 that are extremely unlikely to be chance associations but have significantly boosted 3D velocity ($v/v_{\rm escN}>1$).
3.3 Systematic errors?
@@
As described in Sections 3.1 and 3.2, the reasonable interpretation of the Saglia:2025 sample indicates tension with Newton in favor of MOND-based modified gravity. Considering the profound implication of this (though moderate) conclusion, it is warranted to investigate any possible systematic effects that can bias the gravity inference.
@@
Can this tension be attributed to an unrecognized bias in the Saglia:2025 sample, or known issues in the Gaia data? Considering the tension in the distribution of $\Delta\phi$ (Section 3.1), one may wonder whether the Saglia:2025 sample selection preferred certain phase (e.g. apastron or periastron). This seems unlikely because the Saglia:2025 selection is primarily based on the availability of the HARPS radial velocities from binaries of diverse separations, and it is impossible to know the phase in selecting each binary without an accurate orbit reconstruction (in other words, the sample selection is blind to the phase). Known biases in Gaia DR3 parallaxes are potential concerns (Lindegren:2021). For the Saglia:2025 sample, the maximum R.A. and decl. separations are $0.111^{\circ}$ and $0.086^{\circ}$. On the scale of $0.1^{\circ}$, the RMS amplitude of the smoothed variations in derived bias patterns (Figure 14 of Lindegren:2021) is $7.7\mu\rm{as}$, which is smaller than the reported minimal nominal error of $12\mu\rm{as}$ in the Saglia:2025 sample. Since the known parallax bias is smaller than the actually used errors of parallaxes, it would have not significantly affected the Bayesian results. Nevertheless, the parallax bias highlights the fact that the current data have significant limitations with regard to the radial separation.
@@
The input stellar masses can potentially have the largest systematic effect. Saglia:2025 estimated stellar masses through state-of-the-art stellar evolution modeling based on a wealth of spectroscopic, photometric, and parallax data, and quoted a systematic error of at most 10% at the 95% confidence level. The good agreement with the Newtonian expectation for the 24 binaries in the high-acceleration regime (the right panel of Figure 8) may strengthen the reliability of their stellar masses. Noting this, it seems contrived to suspect that only the eight wide binaries in the low-acceleration regime have systematically biased stellar masses. Nevertheless, if a 10% increase in stellar mass is uniformly applied to all stars in the eight wide binaries, the consolidated value of $\Gamma$ through the MCMC procedure is adjusted only by $\Delta\Gamma\approx-0.02$, meaning that the statistical significance of the anomaly is only weakly reduced. With the 10% increased mass, the Bayesian outputs in Newtonian gravity for binary #24 would give $v_{\rm obs}/v_{\rm escN}=1.073\pm 0.029$, $\log_{10}f_{M}=0.028\pm 0.019$, and $v_{\rm mod}/v_{\rm obs}=0.90\pm 0.03$.
4 Summary and Outlook
@@
The new algorithm is different from the algorithm by Chae:2025 in that all parameters of the former have correct physical meanings, while the latter is mainly designed to measure $\Gamma$ (e.g., the inclination parameter $i$ is not a true angle between the sky plane and the orbital plane in the latter).
A pilot study with the general algorithm is carried out for the Saglia:2025 32 binaries with precise HARPS RVs. Two approaches are considered to test or infer gravity in wide binaries: statistical testing or direct gravity inference. In the statistical testing approach, distributions of the inferred orbit and orientation parameters are statistically tested for the assumed gravity. In this approach, three models (Newtonian, MOND-type toy, and general) of gravity are considered with or without priors. In the gravity inference approach, the logical priors on inclination $i$ and phase $\Delta\phi$ (the true anomaly parameter) are imposed and the strength of gravity is measured through the statistical consolidation of the individual PDFs of $\Gamma$.
@@
  • The results with Newtonian gravity agree overall with the orbit solutions derived by Saglia:2025 except for system #22 (Stars 36OphB and GJ663A) for which Saglia:2025 obtain an unusually large value of $a/s\approx 40$ (with $e=0.25$) while the Bayesian result with priors predicts $a/s\approx 1$ (with $e=0.96_{-0.14}^{+0.02}$) which is much more natural.
  • For Binary #24 (Stars HD189739 and HD189760), the Bayesian modeling in Newtonian gravity returns $\log_{10}f_{M}=0.035\pm 0.019$ (i.e., $\approx 8$% increase in mass), a semi-major axis $a=51.2_{-21.3}^{+52.1}$ kau (much larger than the sky-projected separation $s=5.78$ kau), and a phase close to the periastron ($\Delta\phi=333_{-17}^{+15}$ degrees). Consequently, the Newtonian model predicts a 3D relative velocity $v_{\rm mod}$ inconsistent with the observed velocity $v_{\rm obs}$ with $v_{\rm mod}/v_{\rm obs}=0.87\pm 0.02$ and $v_{\rm obs}/v_{\rm escN}=1.116\pm 0.030$ confirming that $v_{\rm obs}$ exceeds the predicted Newtonian escape velocity as was noticed by Saglia:2025. However, this binary system can be consistent with MOND-type modified gravity.
@@
  • For 8 wide binaries with $g_{\rm N}<10^{-9}$m s-2, the inferred gravity is boosted: $\Gamma=0.134_{-0.036}^{+0.056}$ (thermal prior) or $0.115_{-0.028}^{+0.048}$ (flat prior). These results are $3.7\sigma$ discrepant with Newton, but can be consistent with MOND-type modified gravity within $2\sigma$. The boosted gravity only in the low acceleration regime is consistent with the results from various recent statistical analyses (mostly based on the sky-projected velocities) by two groups (e.g., Chae:2023; Chae:2024a; Chae:2024b; Hernandez:2023; Hernandez:2024a; HernandezKroupa:2025; Chae:2025; Yoon:2025; see also Hernandez:2024review for a discussion of divergent results).
@@
  • If Binary #24 (Stars HD189739 and HD189760) is excluded from the the sample with $g_{\rm N}<10^{-9}$m s-2, the inferred gravity is $\Gamma=0.063_{-0.041}^{+0.065}$ (thermal prior) or $0.051_{-0.039}^{+0.045}$ (flat prior), which favor MOND only marginally/weakly. However, Binary #24 is extremely unlikely to be a chance association or kinematically contaminated based on the currently available observational information.
@@
The pilot study of this work based on the 32 binaries with precise RVs demonstrates the power of the general Bayesian 3D modeling algorithm to measure gravity at low acceleration. While this work was limited by the small sample size, it is expected that $\Gamma$ will be measured with greater and greater precision in the low-acceleration regime based on larger and larger samples of wide binaries with precise RVs. The 8 wide binaries in the low-acceleration regime from Saglia:2025 include one system that is extremely unlikely to be a chance association, but has too large a 3D relative velocity between the pair to be consistent with Newton. It will be interesting to see whether the system happens to be a rare exception/fluke or more such systems are discovered from larger samples of similar data qualities in the future. Also, distributions of the inferred orbit parameters (in particular, the phase true anomaly parameter $\Delta\phi$) in a fixed gravity model may provide a decisive test of the assumed gravity with much larger samples of wide binaries in the future. Thus, standard and nonstandard gravities may be decisively distinguished in the future. Furthermore, the Bayesian 3D modeling methodology may ever more precisely characterize gravity as a function of acceleration to precisely constrain gravity theory beyond standard gravity.
@@
It is the author’s great pleasure to thank the anonymous referee and a Statistics Editor of the AAS journals for useful suggestions including valuable statistical methods that prompted additional analyses and led to much more improved presentations. The author also thanks a data editor for the suggestion of checking and giving the identifications for all the stars of the Saglia:2025 sample using the SIMBAD Astronomical Database (see Table 1 for the correction of two typos in their tables). The author gratefully acknowledges interesting discussions with Xavier Hernandez, Arthur Kosowsky, and David Turnshek. The author also thanks all collaborators (including the mentioned scientists) of follow-up works. The initial results from this work were presented at a PITT PACC Astrophysics Seminar (28 August 2025, University of Pittsburgh), at the workshop “MOND: alternative paths in the dark matter problem” (1 - 5 September 2025) held at the Lorentz center of Leiden University, [^3] https://www.lorentzcenter.nl/mond-alternative-paths-in-the-dark-matter-problem.html and by an invited talk given online at 2025 China-Japan-Korea International Workshop on the Milky Way and Exoplanets (22 - 26 September 2025) held at Hangzhou, China. This work was supported by the National Research Foundation of Korea (NRF-2022R1A2C1092306).

### v5

@@ III.2 Bayesian Inference of the Gravitational Anomaly Parameter $\Gamma$
While Newtonian gravity requires Binary #24 to be a fly-by or chance-association in the 3D space, it is not so, based on MOND gravity theories, because $v/v_{\rm esc}<1$ (see Table 2) is feasible in MOND (existing MOND gravity theories (Bekenstein and Milgrom, 1984; Milgrom, 2010) predict a velocity boost of about 20%). Thus, if the scientific question is to test standard and nonstandard theories with wide binaries, #24 could be taken as evidence for nonstandard theories. It is also important to remember that this system survived stringent observational selection criteria just like all the other systems carefully collected by Saglia et al. (2025). Moreover, the random chance-association probability for this system is extremely low or essentially zero as can be estimated through a Monte Carlo method. Specifically, the expected number of chance-association systems within 100 pc (like #24) for the sky-projected separation <6<6 kau, the line-of-sight separation <0.5<0.5 pc (considering $2\sigma$ uncertainties in the distance measurements), and the 3D relative speed <1<1 km s-1, is $\lesssim 8\times 10^{-4}$ for the one-dimensional velocity dispersion >30>30 km s-1 of peculiar velocities which is a conservative limit for solar-type stars in the solar neighborhood. Moreover, if the scatter of metallicities is considered, the chance gets even lower because two stars of Binary #24 have nearly identical metallicities (Saglia et al., 2025).
Another possibility to consider for Binary #24 is a chance that it became gravitationally unbound just recently through the impulse due to a passing field star. Banik and Zhao (2018) considered this “gravitational ionization” for a binary with a 3D separation of 20 kau and found that the chance is negligibly small, because it will disperse too quickly to be observable during the age of the disk. Thus, it is extremely unlikely that Binary #24 is a gravitationally unbound chance association in 3D space.
@@
Figure 9 explicitly compares the observed 3D velocity components and the total mass of Binary #24 with the best-fit (MLE) values predicted by Newtonian, MOND-type ($G=1.4G_{\rm N}$), and generalized gravity ($\Gamma$ free) models. Newtonian model shows a large discrepancy in $v_{z^{\prime}}$ and a moderate discrepancy in $\log_{10}f_{M}$ while the other models reproduce the observed quantities well. The difference between the Newtonian prediction and the observed value of $v_{z^{\prime}}$ is $111$ m s-1 which is $7.6$ times the nominal measurement uncertainty and $2.6$ times the enlarged uncertainty considering possible errors in correcting for the effects of gravitational redshift and convective flow (see Saglia et al. 2025). Based on the enlarged error of $v_{z^{\prime}}$, $\Delta{\rm BIC}$ (see Appendix A) between the Newtonian model and the other models is $\approx 8$ which is very significant. Moreover, the Newtonian model predicts an extremely unlikely (or impossible) combination of $(e,\Delta\phi)=(1.00,338.2^{\circ})$. It is interesting to note that the pseudo-Newtonian model with $G=1.4G_{\rm N}$ predicts $(e,\Delta\phi)=(0.91,339.7^{\circ})$, which seems somewhat rare but not impossible, while the generalized gravity model predicts $(e,\Delta\phi)=(0.67,203.6^{\circ})$. Because $z^{\prime}$ is not used in obtaining the MLE (see Appendix A), no models match the observed $z^{\prime}$ (nearly) exactly. Interestingly, all models predict $z^{\prime}\approx 0$, which is within its (large) measurement uncertainty.
@@
Although it is extremely unlikely that Binary #24 is a chance association, I also consider results without it just in case the system may be kinematically contaminated (see Section III.3 below.) The thick dotted red curves in the left panels of Figure 8 show the consolidated PDFs of $\Gamma$: $\Gamma=0.063_{-0.041}^{+0.065}$ (thermal prior) or $0.051_{-0.039}^{+0.045}$ (flat prior). These results are only weakly discrepant with Newton and cannot distinguish well between Newton and MOND. This is not surprising given the small number of wide binaries. What is surprising is the presence of Binary #24 whose inclusion results in a discrepancy with Newton. It will be interesting to see whether future observational studies can discover more wide binaries like Binary #24 that are extremely unlikely to be chance associations but have significantly boosted 3D velocity ($v_{\rm obs}/v_{\rm escN}>1$). Finally, it is worth noting that the median value of $\Gamma=0.063$ without Binary #24 means a gravity boost factor of $\gamma\approx 1.3$ that agrees with MOND numerical solutions by Pflamm-Altenburg (2025).
III.3 Systematic errors?
@@
As described in Sections III.1 and III.2, the reasonable interpretation of the Saglia et al. (2025) sample indicates tension with Newton in favor of MOND-based modified gravity. Considering the profound implication of this (though moderate) conclusion, it is warranted to investigate any possible systematic effects that can bias the gravity inference.
@@
Can this tension be attributed to an unrecognized bias in the Saglia et al. (2025) sample, or known issues in the Gaia data? Considering the tension in the distribution of $\Delta\phi$ (Section III.1), one may wonder whether the Saglia et al. (2025) sample selection preferred certain phase (e.g. apastron or periastron). This seems unlikely because the Saglia et al. (2025) selection is primarily based on the availability of the HARPS radial velocities from binaries of diverse separations, and it is impossible to know the phase in selecting each binary without an accurate orbit reconstruction (in other words, the sample selection is blind to the phase). Known biases in Gaia DR3 parallaxes are potential concerns (Lindegren et al., 2021). For the Saglia et al. (2025) sample, the maximum R.A. and decl. separations are $0.111^{\circ}$ and $0.086^{\circ}$. On the scale of $0.1^{\circ}$, the RMS amplitude of the smoothed variations in derived bias patterns (Figure 14 of Lindegren et al. (2021)) is $7.7\mu\rm{as}$, which is smaller than the reported minimal nominal error of $12\mu\rm{as}$ in the Saglia et al. (2025) sample. Since the known parallax bias is smaller than the actually used errors of parallaxes, it would have not significantly affected the Bayesian results. Nevertheless, the parallax bias highlights the fact that the current data have significant limitations with regard to the radial separation.
@@
The input stellar masses can potentially have the largest systematic effect. Saglia et al. (2025) estimated stellar masses through state-of-the-art stellar evolution modeling based on a wealth of spectroscopic, photometric, and parallax data, and quoted a systematic error of at most 10% at the 95% confidence level. The good agreement with the Newtonian expectation for the 24 binaries in the high-acceleration regime (the right panel of Figure 8) may strengthen the reliability of their stellar masses. Noting this, it seems contrived to suspect that only the eight wide binaries in the low-acceleration regime have systematically biased stellar masses. Nevertheless, if a 10% increase in stellar mass is uniformly applied to all stars in the eight wide binaries, the consolidated value of $\Gamma$ through the MCMC procedure is adjusted only by $\Delta\Gamma\approx-0.02$, meaning that the statistical significance of the anomaly is only weakly reduced. With the 10% increased mass, the Bayesian outputs in Newtonian gravity for binary #24 would give $v_{\rm obs}/v_{\rm escN}=1.073\pm 0.029$, $\log_{10}f_{M}=0.028\pm 0.019$, and $v_{\rm mod}/v_{\rm obs}=0.90\pm 0.03$.
Given that the statistical significance of $\Gamma>0$ from this pilot study is largely driven by one system (Binary #24), some further consideration of this system is warranted. Since a chance association is extremely unlikely for this system as discussed above, the remaining possibility for systematic error is kinematic contamination (e.g., Clarke 2020) by an undetected companion within the Gaia angular resolution limit of $\approx 0.5$ - $1.0$ arcsec (Gaia Collaboration et al., 2018), which corresponds to $\approx 50$ - 100 kau for the distance of $\approx 100$ pc. There are several observational tests of kinematic contamination by a hidden close companion, including Gaia’s ruwe value (Belokurov et al., 2020), location in the color-magnitude(CM) diagram, variability of RVs in time (El-Badry, 2024), and direct high-resolution imaging by Speckle interferometry (e.g., Manchanda et al. 2023; Tokovinin 2023).
The two stars HD189739 and HD189760 have fully acceptable ruwe values of $1.103$ and $0.969$ and normal luminosities at their measured Gaia BP-RP colors. The HARPS value of the relative RV $v_{r}=-0.861\pm 0.015$ is well consistent with the Gaia DR3 value of $1.092\pm 0.239$ (although the precision of the latter is relatively poor), indicating stability over $\gtrsim 3$ years. Thus, nominal interpretation would be that both are single stars. However, a companion of 0.1 - 0.25 $M_{\odot}$ and orbit radius of 15 - 30 au, can still pass the ruwe test due to long periods, and being $\lesssim 1$% of the solar luminosity and a fraction of an arcsec separation, will not be obvious in the Gaia DR3 or currently available ground-based imaging or 3-year baseline measurements of RVs. Such an undetected companion can induce a perturbation required for the observed boosted velocity. Thus, an RV monitoring over a longer time baseline and direct Speckle high-resolution imaging are needed to completely resolve the issue. One caveat of this consideration is that from an observational point of view, the same concern can be applicable to all the other 7 systems in the low-acceleration regime, because all satisfy the same selection criteria. For example, if the chance of a hidden companion is 1/8, one has to consider removing any of the 8 systems with an equal probability from the observational point of view.
IV Summary and Outlook
@@
The new algorithm is different from the algorithm by Chae (2025b) in that all parameters of the former have correct physical meanings, while the latter is mainly designed to measure $\Gamma$ (e.g., the inclination parameter $i$ is not a true angle between the sky plane and the orbital plane in the latter).
A pilot study with the general algorithm is carried out for the Saglia et al. (2025) 32 binaries with precise HARPS RVs. Two approaches are considered to test or infer gravity in wide binaries: statistical testing or direct gravity inference. In the statistical testing approach, distributions of the inferred orbit and orientation parameters are statistically tested for the assumed gravity. In this approach, three models (Newtonian, MOND-type toy, and general) of gravity are considered with or without priors. In the gravity inference approach, the logical priors on inclination $i$ and phase $\Delta\phi$ (the true anomaly parameter) are imposed and the strength of gravity is measured through the statistical consolidation of the individual PDFs of $\Gamma$.
@@
  • The results with Newtonian gravity agree overall with the orbit solutions derived by Saglia et al. (2025) except for system #22 (Stars 36OphB and GJ663A) for which Saglia et al. (2025) obtain an unusually large value of $a/s\approx 40$ (with $e=0.25$) while the Bayesian result with priors predicts $a/s\approx 1$ (with $e=0.96_{-0.14}^{+0.02}$) which is much more natural.
  • For Binary #24 (Stars HD189739 and HD189760), [^5] Gaia DR3 identifiers: 4235732073427592704 and 4235731867269159680. the Bayesian modeling in Newtonian gravity returns $\log_{10}f_{M}=0.035\pm 0.019$ (i.e., $\approx 8$% increase in mass), a semi-major axis $a=51.2_{-21.3}^{+52.1}$ kau (much larger than the sky-projected separation $s=5.78$ kau), and a phase close to the periastron ($\Delta\phi=333_{-17}^{+15}$ degrees). Consequently, the Newtonian model predicts a 3D relative velocity $v_{\rm mod}$ inconsistent with the observed velocity $v_{\rm obs}$ with $v_{\rm mod}/v_{\rm obs}=0.87\pm 0.02$ and $v_{\rm obs}/v_{\rm escN}=1.116\pm 0.030$ confirming that $v_{\rm obs}$ exceeds the predicted Newtonian escape velocity as was noticed by Saglia et al. (2025). However, this binary system can be consistent with MOND-type modified gravity.
@@
  • For 8 wide binaries with $g_{\rm N}<10^{-9}$m s-2, the inferred gravity is boosted: $\Gamma=0.134_{-0.036}^{+0.056}$ (thermal prior) or $0.115_{-0.028}^{+0.048}$ (flat prior). For 6 wide binaries with $g_{\rm N}<10^{-9.5}$m s-2, it is found that $\Gamma=0.143_{-0.041}^{+0.068}$ (thermal prior) or $0.149_{-0.044}^{+0.054}$ (flat prior). These results are $\gtrsim 3.5\sigma$ discrepant with Newton, but can be consistent with MOND-type modified gravity within $2\sigma$.
@@
  • If Binary #24 is excluded from the the sample with $g_{\rm N}<10^{-9}$m s-2, the inferred gravity is $\Gamma=0.063_{-0.041}^{+0.065}$ (thermal prior) or $0.051_{-0.039}^{+0.045}$ (flat prior), which would be well consistent with MOND but also consistent with Newton within $1.5\sigma$. For Binary #24, a chance association is extremely unlikely. Also, there is no obvious observational indication for kinematical contamination by a hidden companion. However, possibilities of kinematic contamination cannot be completely ruled out based on the currently available observational information, and thus further observational studies are required.
The above results with all 32 wide binaries from the Saglia et al. (2025) sample are consistent with the results from various recent statistical analyses (mostly based on the sky-projected velocities) by two groups (e.g., Chae 2023, 2024b, 2024a; Hernandez 2023; Hernandez et al. 2024b, a; Hernandez and Kroupa 2025; Chae 2025b; Yoon et al. 2025). However, if Binary #24 is excluded (e.g., assuming that it will turn out to be kinematically contaminated from future observations), the distinction between Newton and MOND is largely removed because of the small sample size. Considering the importance of this system and the recent wide binary results in the literature supporting Newton (e.g., Pittordis and Sutherland 2023; Banik et al. 2024; Pittordis et al. 2025; Makarov 2026), [^6] Since this work is mainly intended to be a presentation of the 3D methodology, a review of recent works is not considered here. it will be interesting to see whether the system happens to be a rare exception/fluke or more such systems are discovered from larger samples of similar data qualities in the future.
@@
The pilot study of this work based on the 32 binaries with precise RVs demonstrates the power of the general Bayesian 3D modeling algorithm to measure gravity at low acceleration. While this work was limited by the small sample size, it is expected that $\Gamma$ will be measured with greater and greater precision in the low-acceleration regime based on larger and larger samples of wide binaries with precise RVs. Also, distributions of the inferred orbit parameters (in particular, the phase true anomaly parameter $\Delta\phi$) in a fixed gravity model may provide a decisive test of the assumed gravity with much larger samples of wide binaries in the future. Thus, standard and nonstandard gravities may be decisively distinguished through 3D velocity data in the future. Larger samples of 3D velocity data than the Saglia et al. (2025) sample are currently being gathered by multiple groups around the world (see, e.g., Saad and Ting 2025). The methodology presented here may provide a useful avenue in interpreting future data for experimental studies of low-acceleration gravity.
@@
It is the author’s great pleasure to thank two anonymous referees and the Statistics Editor of the AAS journals for useful suggestions including valuable statistical methods that prompted additional analyses and led to much more improved presentations. The author also thanks a Data Editor for the suggestion of checking and giving the identifications for all the stars of the Saglia et al. (2025) sample using the SIMBAD Astronomical Database (see Table 1 for the correction of two typos in their tables). The author gratefully acknowledges interesting discussions with Xavier Hernandez, Arthur Kosowsky, Luca Pasquini, and David Turnshek. The author also thanks all collaborators (including the mentioned scientists) of follow-up works. Finally, the author would like to thank the Scientific Editor Fred Rasio for the patient and thoughtful handling of this review process. The initial results from this work were presented at a PITT PACC Astrophysics Seminar (28 August 2025, University of Pittsburgh), at the workshop “MOND: alternative paths in the dark matter problem” (1 - 5 September 2025) held at the Lorentz center of Leiden University, [^7] https://www.lorentzcenter.nl/mond-alternative-paths-in-the-dark-matter-problem.html and by an invited talk given online at 2025 China-Japan-Korea International Workshop on the Milky Way and Exoplanets (22 - 26 September 2025) held at Hangzhou, China. This work was supported by the National Research Foundation of Korea (NRF-2022R1A2C1092306).